# OOTDiffusion: Outfitting Fusion based Latent Diffusion for Controllable Virtual Try-on

### v1

@@ Abstract
Image-based virtual try-on (VTON), which aims to generate an outfitted image of a target human wearing an in-shop garment, is a challenging image-synthesis task calling for not only high fidelity of the outfitted human but also full preservation of garment details. To tackle this issue, we propose Outfitting over Try-on Diffusion (OOTDiffusion), leveraging the power of pretrained latent diffusion models and designing a novel network architecture for realistic and controllable virtual try-on. Without an explicit warping process, we propose an outfitting UNet to learn the garment detail features, and merge them with the target human body via our proposed outfitting fusion in the denoising process of diffusion models. In order to further enhance the controllability of our outfitting UNet, we introduce outfitting dropout to the training process, which enables us to adjust the strength of garment features through classifier-free guidance. Our comprehensive experiments on the VITON-HD and Dress Code datasets demonstrate that OOTDiffusion efficiently generates high-quality outfitted images for arbitrary human and garment images, which outperforms other VTON methods in both fidelity and controllability, indicating an impressive breakthrough in virtual try-on. Our source code is available at https://github.com/levihsu/OOTDiffusion.
Keywords:
Virtual try-on Latent diffusion Outfitting fusion
@@ 1 Introduction
Image-based virtual try-on (VTON) is a popular and promising image-synthesis technology for the current e-commerce industry, which is able to dramatically improve the shopping experience of consumers and reduce the advertising cost of clothing merchants. As its name suggests, the objective of the VTON task is to generate an image of a target human wearing a given garment, which took tremendous efforts from numerous researchers [16, 49, 15, 8, 30, 54, 12, 11, 17, 52, 24] for more natural and accurate virtual try-on results in the past few years.
@@
Image-based VTON is currently facing two main challenges. First, the generated images should be realistic and natural enough to avoid dissonance. Most of recent researches on virtual try-on leverage generative adversarial networks [13] (GANs) or latent diffusion models [40] (LDMs) for image generation. Previous GAN-based methods [16, 15, 6, 17, 27, 52] usually have difficulty in generating correct garment folds, natural light and shadow, and realistic human body, and are prone to overfitting. Hence more recent work favors LDM-based methods [60, 32, 14, 24], which effectively enhance the fidelity of outfitted images. The second critical challenge is how to preserve as much as possible the detail features of garments, such as complicated text, textures, colors, patterns and lines, etc. Previous researches [6, 27, 32, 52, 14] usually perform an explicit warping process to align the garment features with the target human body, and then feed the warped garment into generative models (i.e., GANs and LDMs, etc.). Thus the performance of such VTON methods is extremely dependent on the efficacy of the lossy warping process. On the other hand, some LDM-based methods [32, 14, 24] attempt to learn garment features via CLIP textual-inversion [10], which fail to preserve fine-grained garment details.
@@
Motivated by the aforementioned prospects and challenges of image-based VTON, we propose a novel LDM-based virtual try-on method, namely Outfitting over Try-on Diffusion (OOTDiffusion; see Fig. 2). First, we make full use of the advantages of pretrained latent diffusion models [40] to ensure high fidelity of generated images and natural try-on effects. Then we introduce a novel outfitting UNet to learn the precise garment features in the latent space, and merge them with the noisy human body in the denoising UNet through an outfitting fusion process in the self-attention layers [48]. In this way, the garment features are smoothly adapted to various target human body types and postures, without suffering information loss or feature distortion caused by warping. Meanwhile, we also conduct CLIP textual-inversion [10] for the garment image as the auxiliary semantic information fed into both outfitting and denoising UNets via the cross-attention mechanism [48]. Furthermore, we perform an outfitting dropout process, which randomly drops a handful of garment features in training to enable classifier-free guidance [20] with respect to the garment features. Through this approach, the strength of garment control over the generated result can be simply adjusted by a guidance scale, which further enhances the controllability of our VTON method.
@@
  • We propose OOTDiffusion, a novel LDM-based network architecture for realistic and controllable virtual try-on.
  • We design an outfitting UNet to learn the precise garment features and align them with the target human body via our outfitting fusion in the self-attention layers without an explicit warping process.
  • We introduce outfitting dropout to the training process, which further improves the controllability of our outfitting UNet.
  • We train our OOTDiffusion on two broadly-used high-resolution benchmark datasets, i.e., VITON-HD [6] and Dress Code [33], respectively. Extensive qualitative and quantitative evaluations demonstrate our superiority over state-of-the-art VTON methods in both fidelity and controllability for various target human and garment images (see Fig. 1), which implies an impressive breakthrough in image-based virtual try-on.
@@ 2 Related Work
Image-based virtual try-on has been investigated for many years as a promising and challenging task [60, 49, 22, 16, 9, 14, 33, 52, 24, 27, 32, 6]. Aiming at more natural and accurate results, recent researches are mainly based on generative adversarial networks [13] (GANs) or latent diffusion models [40] (LDMs) for image generation. Among the GAN-based VTON methods [6, 27, 52], VITON-HD [6] collected a high-resolution dataset and proposed ALIAS normalization and generator to address the misalignment between warped clothes and target regions. HR-VITON [27] simultaneously performed warping and segmentation to handle the body occlusion and garment misalignment. GP-VTON [52] proposed a LFGP warping module to generate deformed garments and introduced a DGT training strategy for the warping network. As introduced above, GAN-based methods usually rely on an explicit warping process neglecting realistic garment folds and natural light and shadow, which seriously degrades the fidelity and realism of outfitted images. Meanwhile, GAN-based methods are prune to overfitting and causing severe performance degradation on out-of-distribution images.
With respect to the LDM-based approaches [32, 14, 24], LaDI-VTON [32] and DCI-VTON [14] also require an explicit warping process. In specific, LaDI-VTON [32] performed textual-inversion to map the visual garment features to the CLIP [39] token embedding space and condition the latent diffusion model along with the warped input. DCI-VTON [14] directly combined the warped clothes with the masked person image to get a coarse result, and then refined it by the diffusion model. Neither of these methods succeeded in fully preserving garment details like complicated pattern and text due to the information loss caused by the CLIP encoder. More recently, StableVITON [24] discarded independent warping and proposed a zero cross-attention block to learn semantic correlation between the clothes and human body. However, information loss still exists in cross-attention and the extra zero-initialized blocks heavily increase its training and inference cost. In contrast, again without warping, our LDM-based OOTDiffusion finetunes the pretrained outfitting UNet to learn garment details in one step and efficiently merges them into the denoising UNet via our outfitting fusion with negligible information loss.
@@ 3.2 OOTDiffusion
Fig. 2 illustrates the overview of our method. Given a target human image $\mathbf{x}\in\mathbb{R}^{3\times H\times W}$ and an input garment image $\mathbf{g}\in\mathbb{R}^{3\times H\times W}$, our OOTDiffusion is capable of generating a realistic outfitted image $\mathbf{x_{g}}\in\mathbb{R}^{3\times H\times W}$. We employ OpenPose [4, 46, 5, 51] and HumanParsing [28] to generate a masked human image $\mathbf{x_{m}}\in\mathbb{R}^{3\times H\times W}$, and use a VAE encoder $\mathcal{E}$ to transform it into the latent space as $\mathcal{E}(\mathbf{x_{m}})\in\mathbb{R}^{4\times h\times w}$, where $h=\frac{H}{8}$ and $w=\frac{W}{8}$. Then we concatenate $\mathcal{E}(\mathbf{x_{m}})$ with a Gaussian noise $\epsilon\in\mathbb{R}^{4\times h\times w}$ as the input latent $\mathbf{z}_{T}\in\mathbb{R}^{8\times h\times w}$ for the denoising UNet. Note that we add $4$ zero-initialized channels to the first convolutional layer of the denoising UNet to support our input with $8$ channels.
@@
As introduced above, we propose an outfitting UNet to efficiently learn the detail features of the garment image $\mathbf{g}$. The left side of Fig. 2 shows the architecture of our outfitting UNet, which is essentially identical to the denoising UNet of Stable Diffusion. The encoded garment latent $\mathcal{E}(\mathbf{g})\in\mathbb{R}^{4\times h\times w}$ is fed into the outfitting UNet $\omega_{\theta^{\prime}}$, and then incoporated into the denoising UNet $\epsilon_{\theta}$ via our outfitting fusion (see the next section). Along with the aforementioned auxiliary conditioning input, the outfitting and denoising UNets are jointly trained by minimizing the following loss function:
@@
where $\psi=\tau_{g}(\mathbf{g})\ \textcircled{c}\ \tau_{y}(\mathbf{y})$ represents the auxiliary conditioning input for both $\omega_{\theta^{\prime}}$ and $\epsilon_{\theta}$. While $\tau_{g}$ and $\tau_{y}$ refer to the pretrained CLIP image encoder and text encoder respectively, and c⃝ denotes concatenation.
@@
Based on our proposed outfitting UNet and inspired by the spatial-attention mechanism [48, 21], we introduce an outfitting fusion process to incorporate the learned garment features into the denoising UNet. First, we dive into the transformer blocks [48] of two UNets, finding each pair of feature maps used as input to the corresponding self-attention layers [48]. Given the $n$th pair of feature maps $\mathbf{g}_{n},\mathbf{x}_{n}\in\mathbb{R}^{c_{n}\times h_{n}\times w_{n}}$, we concatenate them in the spatial domain as:
$\mathbf{x_{g}}_{n}=\mathbf{x}_{n}\ \textcircled{c}\ \mathbf{g}_{n}\in\mathbb{R}^{c_{n}\times h_{n}\times 2w_{n}}.$ (3)
Then we replace $\mathbf{x}_{n}$ with the concatenated feature map $\mathbf{x_{g}}_{n}$ as the input to the self-attention layer of the denoising UNet. After the self-attention process, we crop out the fist half of the feature map as the output of the self-attention layer. Through this approach, the garment features are effectively correlated with the noisy human body with negligible information loss. Hence the denoising UNet is made capable of learning the precise features from the outfitting UNet for preserving garment details and naturally adapting them to the target human body in the generated image.
@@
In order to further enhance the controllability of our VTON method, we employ an outfitting dropout operation in training to enable classifier-free guidance [20] for the garment features. Classifier-free guidance has been broadly used in conditional image generation [35, 56, 44, 3] for trading off the quality and diversity of images generated by latent diffusion models. Specifically in the training process of our outfitting UNet, we randomly drop the input garment latent as $\mathcal{E}(\mathbf{g})=\varnothing$, where $\varnothing\in\mathbb{R}^{4\times h\times w}$ refers to an all-zero latent. In this way, the denoising UNet is trained both conditionally and unconditionally, i.e., with and without the outfitting fusion. Then at inference time, we simply use a guidance scale $s_{\mathbf{g}}\geq 1$ to adjust the strength of conditional control over the predicted noise $\hat{\epsilon}_{\theta}$ as:
@@
where we omit some minor terms compared with Eq. 2 for the sake of brevity.
In practice, we empirically set the outfitting dropout ratio to $10\%$ in training, i.e., $10\%$ of garment latents $\mathcal{E}(\mathbf{g})$ are set to $\varnothing$. And the optimal value of the guidance scale $s_{\mathbf{g}}$ is usually around $1.5\sim 2.0$ according to our ablation study (see Sec. 4.3). Fig. 3 and Tab. 1 demonstrate the effects of our outfitting dropout and different guidance scale values.
@@ 4.1 Experimental Setup
We evaluate the results in both the paired and unpaired settings, where the paired setting provides the target human and the corresponding garment images for reconstruction, and the unpaired setting provides the different garment images for virtual try-on. Specifically for Dress Code [33], we note that the evaluation are performed on the entire dataset rather than being limited to upper-body garments, which more effectively validates the feasibility of each method in real-world applications with various garment types. In addition, we conduct cross-dataset evaluation, in which the models are trained on one dataset and evaluated on the other, to further investigate the generalization ability of each method.
@@ 4.2 Implementation Details
In our experiments, we initialize the OOTDiffusion models by inheriting the pretrained weights of Stable Diffusion v1.5 [40]. Then we finetune the outfitting and denoising UNets using an AdamW optimizer [29] with a fixed learning rate of 5e-5. Note that we train four types of models on VITON-HD [6] and Dress Code [33] datasets at resolutions of $512\times 384$ and $1024\times 768$, separately. All the models are trained for 36000 iterations on a single NVIDIA A100 GPU, with a batch size of 64 for the $512\times 384$ resolution and 16 for the $1024\times 768$ resolution. At inference time, we run our OOTDiffusion on a single NVIDIA RTX 4090 GPU for 20 sampling steps using the UniPC sampler [59].
@@ 4.3 Ablation Study
Table 1: Ablation study of outfitting dropout and different guidance scale values on the VITON-HD dataset. The best and second best results are reported in bold and underline, respectively.
@@
We investigate the effects of our proposed outfitting dropout as well as the different values of the guidance scale $s_{\mathbf{g}}$ on the VITON-HD dataset [6]. First, we train two variants of our OOTDiffusion models without/with outfitting dropout, respectively. Then for the model trained with outfitting dropout, we set $s_{\mathbf{g}}=1.0,1.5,2.0,2.5,3.0,5.0$ for classifier-free guidance. At inference time, we guarantee all of other parameters (including the random seed) are consistent for fair comparison. As Fig. 3 shows, without outfitting dropout, classifier-free guidance is not supported and the generated result is obviously the worst. While for the model trained with outfitting dropout, when $s_{\mathbf{g}}=1.0$, the inference process is identical to the model without outfitting dropout (see Eq. 4), which gets a similarly bad result. When $s_{\mathbf{g}}>1.0$, we see that the fine-grained garment features become clearer as $s_{\mathbf{g}}$ increases. However, color distortion occurs when $s_{\mathbf{g}}\geq 2.5$ and becomes extremely significant when $s_{\mathbf{g}}=5.0$. Furthermore, Tab. 1 quantitatively proves the efficacy of our outfitting dropout which enables classifier-free guidance with respect to the garment features, and find the optimal guidance scale value is around $1.5\sim 2.0$ in most cases. According to this study, we consistently conduct outfitting dropout for OOTDiffusion, and empirically set $s_{\mathbf{g}}=1.5$ for the VITON-HD dataset [6] and $s_{\mathbf{g}}=2.0$ for the Dress Code dataset [33] in the following experiments.
4.4 Experimental Results
4.4.1 Qualitative Results.
@@ 4.4 Experimental Results
Fig. 4 visually shows some example results of our method and other VTON methods on the test set of VITON-HD [6]. We observe that compared with other methods, our OOTDiffusion consistently achieves the best try-on effect for various upper-body garments. More specifically, GAN-based methods like GP-VTON [52] often fail to generate realistic human bodies (1st and 4th rows) and natural garment folds (2nd and 3rd rows), which makes the outfitted images look unrealistic. While other LDM-based methods like LaDI-VTON [32] and StableVITON [24] tend to lose some garment details such as complicated text (2nd and 4th rows) and patterns (1st and 3rd rows). In contrast, our OOTDiffusion not only generates realistic images but also preserves most of the garment details.
Regarding the more complicated Dress Code dataset [33], which consists of full-body models and various garment categories, our OOTDiffusion still visually outperforms other VTON methods. As illustrated in Fig. 5, Paint-by-Example [53] and LaDI-VTON [32] fail to preserve the garment features, and GP-VTON [52] tends to cause severe body and background distortion. On the contrary, our OOTDiffusion consistently shows very stable performance on different garment categories including upper-body garments (1st row), lower-body garments (2nd row) and dresses (3rd and 4th rows).
In order to evaluate the generalization ability of our method, we conduct an additional cross-dataset experiment, i.e., training on one dataset and testing on the other. Fig. 6 demonstrates that among all the models trained on the VITON-HD dataset [6], our OOTDiffusion is optimally adapted to the test examples in the Dress Code dataset [33], generating more realistic outfitted images and preserving much more garment details. In summary, the observations above (Figs. 4, 5 and 6) qualitatively prove the superiority and generalization capability of our OOTDiffusion in generating natural and accurate outfitted results for various human and garment images.

### v2

@@ Abstract
We present OOTDiffusion, a novel network architecture for realistic and controllable image-based virtual try-on (VTON). We leverage the power of pretrained latent diffusion models, designing an outfitting UNet to learn the garment detail features. Without a redundant warping process, the garment features are precisely aligned with the target human body via the proposed outfitting fusion in the self-attention layers of the denoising UNet. In order to further enhance the controllability, we introduce outfitting dropout to the training process, which enables us to adjust the strength of the garment features through classifier-free guidance. Our comprehensive experiments on the VITON-HD and Dress Code datasets demonstrate that OOTDiffusion efficiently generates high-quality try-on results for arbitrary human and garment images, which outperforms other VTON methods in both realism and controllability, indicating an impressive breakthrough in virtual try-on. Our source code is available at https://github.com/levihsu/OOTDiffusion.
@@ 1 Introduction
Image-based virtual try-on (VTON) is a popular and promising image-synthesis technology for the current e-commerce industry, which is able to dramatically improve the shopping experience of consumers and reduce the advertising cost of clothing merchants. As its name suggests, the VTON task aims to generate an outfitted image of a target human wearing a given garment. It has taken tremendous efforts from numerous researchers [16, 49, 15, 8, 30, 54, 12, 11, 17, 52, 24] for more natural and accurate virtual try-on results in the past few years.
@@
Image-based VTON is currently facing two main challenges. First, the generated images should be realistic and natural enough to avoid dissonance. Most of recent researches on virtual try-on leverage generative adversarial networks [13] (GANs) or latent diffusion models [40] (LDMs) for image generation. Previous GAN-based methods [16, 15, 6, 17, 27, 52] usually have difficulty in generating correct garment folds, natural light and shadow, or realistic human bodies. Hence more recent work favors LDM-based methods [60, 32, 14, 24], which effectively improve the realism of outfitted images. The second critical challenge is how to preserve as much as possible the garment detail features, such as complicated text, textures, colors, patterns and lines, etc. Previous researches [6, 27, 32, 52, 14] usually perform an explicit warping process to align the garment features with the target human body, and then feed the warped garment into generative models (i.e., GANs and LDMs, etc.). Thus the performance of such VTON methods is extremely dependent on the efficacy of the independent warping process which is prone to overfitting the training data. On the other hand, some LDM-based methods [32, 14, 24] attempt to learn garment features via CLIP textual-inversion [10], which fail to preserve fine-grained garment details.
@@
Motivated by the aforementioned prospects and challenges of image-based VTON, we present a novel LDM-based virtual try-on method, namely Outfitting over Try-on Diffusion (OOTDiffusion; see Fig.˜2). First, we make full use of the advantages of pretrained latent diffusion models [40] to ensure high realism of generated images and natural try-on effects, and design an outfitting UNet to learn the detail features of garments in the latent space in a single step. Then, we propose an outfitting fusion process to precisely align the garment features with the noisy human body in the self-attention layers [48] of the denoising UNet. In this way, the garment features are smoothly adapted to various target human body types and postures, without suffering information loss or feature distortion caused by an independent warping process. Furthermore, we perform an outfitting dropout operation, randomly dropping a handful of garment latents in training to enable classifier-free guidance [20] with respect to the garment features. Through this approach, the strength of garment control over the generated result can be simply adjusted by a guidance scale, which further enhances the controllability of our VTON method.
@@
  • We present OOTDiffusion, an LDM-based network architecture with a novel outfitting UNet for realistic and controllable virtual try-on.
  • We propose outfitting fusion to efficiently align the garment features with the target human body in the self-attention layers without redundant warping.
  • We introduce outfitting dropout to the training process, which further improves the controllability of the outfitting UNet.
  • We train our OOTDiffusion on two broadly-used high-resolution benchmark datasets, i.e., VITON-HD [6] and Dress Code [33], respectively. Extensive qualitative and quantitative evaluations demonstrate our superiority over the state-of-the-art VTON methods in both realism and controllability for various target human and garment images (see Fig.˜1), implying an impressive breakthrough in image-based virtual try-on.
@@ 2 Related Work
Image-based virtual try-on has been investigated for many years as a promising and challenging task [60, 49, 22, 16, 9, 14, 33, 52, 24, 27, 32, 6]. Aiming at more natural and accurate results, recent researches are mainly based on generative adversarial networks [13] (GANs) or latent diffusion models [40] (LDMs) for image generation. Among the GAN-based VTON methods [6, 27, 52], VITON-HD [6] collected a high-resolution dataset and proposed ALIAS normalization and generator to address the misalignment between warped clothes and target regions. HR-VITON [27] simultaneously performed warping and segmentation to handle the body occlusion and garment misalignment. GP-VTON [52] proposed an LFGP warping module to generate deformed garments and introduced a DGT training strategy for the warping network. As introduced above, GAN-based methods usually rely on an explicit warping process neglecting realistic garment folds and natural light and shadow, which seriously degrades the fidelity and realism of outfitted images. Meanwhile, GAN-based methods are prone to overfitting the training data and causing severe performance degradation on out-of-distribution images.
With respect to the LDM-based approaches [32, 14, 24], LaDI-VTON [32] and DCI-VTON [14] also require an explicit warping process. In specific, LaDI-VTON [32] performed textual-inversion to map the visual garment features to the CLIP [39] token embedding space and condition the latent diffusion model along with the warped input. DCI-VTON [14] directly combined the warped clothes with the masked person image to get a coarse result, and then refined it by the diffusion model. Neither of these methods succeeded in fully preserving garment details like complicated patterns and text due to the information loss caused by the CLIP encoder. More recently, StableVITON [24] discarded independent warping and proposed a zero cross-attention block to learn semantic correlation between the clothes and human body. However, information loss remains in the cross-attention layers, and the extra zero-initialized blocks heavily increase the training and inference cost. In contrast, again without warping, our LDM-based OOTDiffusion finetunes the pretrained outfitting UNet to learn garment details in one step and efficiently incorporates them into the denoising UNet via our outfitting fusion with negligible information loss.
@@ 3.2 OOTDiffusion
Fig.˜2 illustrates the overview of our method. Given a target human image $\mathbf{x}\in\mathbb{R}^{3\times H\times W}$ and an input garment image $\mathbf{g}\in\mathbb{R}^{3\times H\times W}$, OOTDiffusion is capable of generating a realistic outfitted image $\mathbf{x_{g}}\in\mathbb{R}^{3\times H\times W}$. We employ OpenPose [4, 46, 5, 51] and HumanParsing [28] to generate a masked human image $\mathbf{x_{m}}\in\mathbb{R}^{3\times H\times W}$, and use a VAE encoder $\mathcal{E}$ to transform it into the latent space as $\mathcal{E}(\mathbf{x_{m}})\in\mathbb{R}^{4\times h\times w}$, where $h=\frac{H}{8}$ and $w=\frac{W}{8}$. Then we concatenate $\mathcal{E}(\mathbf{x_{m}})$ with a Gaussian noise $\epsilon\in\mathbb{R}^{4\times h\times w}$ as the input latent $\mathbf{z}_{T}\in\mathbb{R}^{8\times h\times w}$ for the denoising UNet. Note that we add $4$ zero-initialized channels to the first convolutional layer of the denoising UNet to support our input with $8$ channels.
@@
As introduced above, we propose an outfitting UNet to efficiently learn the detail features of the garment image $\mathbf{g}$. The left side of Fig.˜2 shows the architecture of our outfitting UNet, which is essentially identical to the denoising UNet of Stable Diffusion. The encoded garment latent $\mathcal{E}(\mathbf{g})\in\mathbb{R}^{4\times h\times w}$ is fed into the outfitting UNet $\omega_{\theta^{\prime}}$, and then incoporated into the denoising UNet $\epsilon_{\theta}$ via our outfitting fusion (see the next section). Along with the aforementioned auxiliary conditioning input, the outfitting and denoising UNets are jointly trained by minimizing the following loss function:
@@
[…]
@@
Figure 3: Visualization of the attention maps with respect to the human body (1st row) and garment features (2nd row) aligned by our outfitting fusion.
Based on our proposed outfitting UNet and inspired by the spatial-attention mechanism [48, 21], we propose an outfitting fusion process to incorporate the learned garment features into the denoising UNet. First, we dive into the transformer blocks [48] of two UNets, finding each pair of feature maps used as the input to the corresponding self-attention layers [48]. Given the $n$th pair of the feature maps $\mathbf{g}_{n},\mathbf{x}_{n}\in\mathbb{R}^{c_{n}\times h_{n}\times w_{n}}$, we concatenate them in the spatial domain as:
[…]
And we replace $\mathbf{x}_{n}$ with the concatenated feature map $\mathbf{x_{g}}_{n}$ as the input to the self-attention layer of the denoising UNet. Then we crop out the fist half of the output feature map as the final output of the self-attention layer. Fig.˜3 visualizes the attention maps learned in our modified self-attention. We observe that the unmasked region focuses attention on the human body (1st row), and the masked pixels are attracted to the garment features (2nd row). Meanwhile, during the denoising process, the attention to the human body gradually includes part of the masked region like the neck and arms, and the attention to the garment features gradually increases in the region of the complicated patterns. Through outfitting fusion in the self-attention layers, the garment features are implicitly warped and effectively correlated with the target human body with negligible information loss. Hence the denoising UNet is made capable of learning the precise features from the outfitting UNet for preserving garment details and naturally adapting them to the target human body in the generated image.
@@
In order to further enhance the controllability of our VTON method, we employ an outfitting dropout operation in training to enable classifier-free guidance [20] with respect to the garment features. Classifier-free guidance has been broadly used in conditional image generation [35, 56, 44, 3] for trading off the quality and diversity of images generated by latent diffusion models. Specifically in the training process of our outfitting UNet, we randomly drop the input garment latent as $\mathcal{E}(\mathbf{g})=\varnothing$, where $\varnothing\in\mathbb{R}^{4\times h\times w}$ refers to an all-zero latent. In this way, the denoising UNet is trained both conditionally and unconditionally, i.e., with and without the outfitting fusion. Then at inference time, we simply use a guidance scale $s_{\mathbf{g}}\geq 1$ to adjust the strength of conditional control over the predicted noise $\hat{\epsilon}_{\theta}$ as:
@@
where we omit some minor terms compared with Eq.˜2 for the sake of brevity.
In practice, we empirically set the outfitting dropout ratio to $10\%$ in training, i.e., $10\%$ of garment latents $\mathcal{E}(\mathbf{g})$ are set to $\varnothing$. And the optimal value of the guidance scale $s_{\mathbf{g}}$ is usually around $1.5\sim 2.0$ according to our ablation study (see Sec.˜4.3). Fig.˜4 and Tab.˜1 demonstrate the effects of our outfitting dropout and different guidance scale values.
@@ 4.1 Experimental Setup
We evaluate the results in both the paired and unpaired settings, where the paired setting provides the target human and the corresponding garment images for reconstruction, and the unpaired setting provides the different garment images for virtual try-on. Specifically for Dress Code [33], we note that the evaluation is performed on the entire dataset rather than being limited to upper-body garments. This more effectively validates the feasibility of each method in real-world applications with various garment types.
@@ 4.2 Implementation Details
In our experiments, we initialize the OOTDiffusion models by inheriting the pretrained weights of Stable Diffusion v1.5 [40]. Then we finetune the outfitting and denoising UNets using an AdamW optimizer [29] with a fixed learning rate of 5e-5. Note that we train four types of models on VITON-HD [6] and Dress Code [33] datasets at resolutions of $512\times 384$ and $1024\times 768$, separately. All the models are trained for 36,000 iterations on a single NVIDIA A100 GPU, with a batch size of 64 for the $512\times 384$ resolution and 16 for the $1024\times 768$ resolution. At inference time, we run our OOTDiffusion on a single NVIDIA RTX 4090 GPU for 20 sampling steps using the UniPC sampler [59].
@@ 4.3 Ablation Study
Table 1: Ablation study of outfitting dropout and different guidance scale values on the VITON-HD dataset [6]. The best and second best results are reported in bold and underline, respectively.
@@
We investigate the effects of our proposed outfitting dropout as well as the different values of the guidance scale $s_{\mathbf{g}}$ on the VITON-HD dataset [6]. First, we train two variants of our OOTDiffusion models without/with outfitting dropout, respectively. Then for the model trained with outfitting dropout, we set $s_{\mathbf{g}}=1.0,1.5,2.0,2.5,3.0,5.0$ for classifier-free guidance. At inference time, we guarantee all of other parameters (including the random seed) are consistent for fair comparison. As Fig.˜4 shows, without outfitting dropout, classifier-free guidance is not supported and the generated result is obviously the worst. While for the model trained with outfitting dropout, when $s_{\mathbf{g}}=1.0$, the inference process is identical to the model without outfitting dropout (see Eq.˜4), which gets a similarly bad result. When $s_{\mathbf{g}}>1.0$, we see that the fine-grained garment features become clearer as $s_{\mathbf{g}}$ increases. However, color distortion occurs when $s_{\mathbf{g}}\geq 2.5$ and becomes extremely significant when $s_{\mathbf{g}}=5.0$ (see the flower patterns). Furthermore, Tab.˜1 quantitatively proves the efficacy of our outfitting dropout which enables classifier-free guidance with respect to the garment features, and finds the optimal guidance scale value is around $1.5\sim 2.0$ in most cases. According to this study, we consistently conduct outfitting dropout for OOTDiffusion, and empirically set $s_{\mathbf{g}}=1.5$ for the VITON-HD dataset [6] and $s_{\mathbf{g}}=2.0$ for the Dress Code dataset [33] in the following experiments.
@@
4.4 Experimental Results
4.4.1 Qualitative Results.
Fig.˜5 visually shows some example results of our method and other VTON methods on the test set of VITON-HD [6]. We observe that compared with other methods, our OOTDiffusion consistently achieves the best try-on effects for various upper-body garments. More specifically, GAN-based methods like GP-VTON [52] often fail to generate realistic human bodies (1st and 4th rows) or natural garment folds (2nd and 3rd rows), making the outfitted images look unrealistic. While other LDM-based methods including LaDI-VTON [32] and StableVITON [24] tend to lose some garment details such as complicated text (2nd and 4th rows) or patterns (1st and 3rd rows). In contrast, our OOTDiffusion not only generates realistic images but also preserves most of the fine-grained garment details.
Regarding the more complicated Dress Code dataset [33], which consists of full-body models and various garment categories, our OOTDiffusion still visually outperforms other VTON methods. As illustrated in Fig.˜6, Paint-by-Example [53] and LaDI-VTON [32] fail to preserve the garment features, and GP-VTON [52] tends to cause severe body and background distortion. On the contrary, our OOTDiffusion consistently shows very stable performance on different garment categories including upper-body garments (1st row), lower-body garments (2nd row) and dresses (3rd and 4th rows).
In order to evaluate the generalization ability of our method, we conduct an additional cross-dataset experiment, i.e., training on one dataset and testing on the other. Fig.˜7 demonstrates that among all the models trained on the VITON-HD dataset [6], our OOTDiffusion is optimally adapted to the test examples in the Dress Code dataset [33], generating more realistic outfitted images and preserving much more garment details. In summary, the observations above (Figs.˜5, 6 and 7) qualitatively prove the superiority and generalization capability of our OOTDiffusion in generating natural and accurate try-on results for various human and garment images.